OpenStudy (anonymous):

For the function f(x)= (x+2)/(x^2+8x+12) name the following : Domain, Holes, VA, HA

OpenStudy (anonymous):

Start by factoring the top and bottom. Look at factors in the bottom, and for each one, determine which x value will make that x 0. Why? Because if that factor is 0, then the whole denominator will be 0, which makes your function undefined. =( Those x values will be missing from your domain. Now, for each of those factors, ask yourself, "Is this factor also a factor of the top?" If it is, then you have a hole at that x. If it is not, then there will be a vertical asymptote at that x.

OpenStudy (cwrw238):

further help: f(x) = (x+2)/(x^2+8x+12) = (x+2) / (x+2)(x+ 6)